OpenStudy (anonymous):

Andrew thinks he has a good thing going. He invests $8,000 with the firm Dewey, Cheatham, and Howe, purchasing a CD (certificate of deposit) that after 10 years will be worth $9,600. The trouble is, Andrew did not consult his investment advisor before putting his money down, and this sneaky company has locked him into an investment that pays simple interest, not compound interest, on his principal. What percent is the company's yearly interest rate? Round to the nearest whole number and do not include units in your answer.

OpenStudy (anonymous):

Final amount=9600 What he invested=8000 Interest earned=1600 rate=a Interest formula 8000*a=1600

OpenStudy (anonymous):

the 't' is missing above post 8000*a*t=1600 8000*a*10=1600

OpenStudy (anonymous):

80000a=1600 a=.02 remember this is decimal, what do you do to get percent?

OpenStudy (anonymous):

move the point 2 untis to the right i shouild really learn to b patient cuz wen u said 9600 i thought that was the answer but then u just gave me the explanation and everything and i got it wrong cuz i was impatient

OpenStudy (anonymous):

So 2% right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

good